OpenStudy (anonymous):

Enrollment at a local university is 8,500. There are one third as many students over the age of 25 as there are 25 and under. How many students are 25 and under and how many are over 25?

OpenStudy (anonymous):

Let x = students 25 and under Then students over 25 = (1/3)x Total students = 8500 and a person is either over 25 or 25 and under, so: (1/3)x + x = 8500 Coefficient of x (1) is same as 3/3, add. (4/3)x = 8500 Times 3/4. x = 6375 students over 25 = (1/3)x Plugging x in, students over 25 = 2125 2125 students > 25, 6375 <= 25